\begin{thebibliography}{100}

\bibitem{Rogo-9} %{bib:01}
W. Rogosinski, “\"{U}ber Bildschranken bei Potenzreihen und ihren Abschnitten,”
Math. Z. \textbf{17}, 260--276 (1923).

\bibitem{LanDai-86} %{bib:02}
E. Landau and D. Gaier,
“Darstellung und Begr\"{u}undung einiger neuerer Ergebnisse der Funktionentheorie,” Springer-Verlag Berlin Heidelberg, 201 (1986).

\bibitem{SchSz-25} %{bib:03}
I. Schur und G. Szeg\"{o},
“\"{U}ber die Abschnitte einer im Einheitskreise beschr\"{a}nkten Potenzreihe,” Sitz.-Ber. Preuss. Acad. Wiss. Berlin Phys.-Math. Kl., 545--560 (1925).

\bibitem{Bohr-14} %{bib:05}
H. Bohr, “A theorem concerning power series,” Proc. London Math. Soc. \textbf{13} (2), 1--5 (1914).

\bibitem{bib:22} C. B\'en\'eteau, A. Dahlner, and D. Khavinson,
“Remarks on the Bohr phenomenon,” Comput. Methods Funct. Theory
\textbf{4} (1), 1--19 (2004).

\bibitem{PPS2002} %{bib:06}
V. I. Paulsen, G. Popescu and D. Singh, “On Bohr’s inequality,”
Proc. London Math. Soc. \textbf{85}, 493--512 (2002).

\bibitem{bib:09} V. I. Paulsen and D. Singh,
“Bohr’s inequality for uniform algebras,” Proc. Amer. Math. Soc. \textbf{132} (12), 3577--3579 (2004).

\bibitem{bib:18} I. R. Kayumov and S. Ponnusamy, “Bohr's inequalities for the analytic functions with lacunary series and harmonic functions,” J. Math. Anal. and Appl., \textbf{465}, 857--871 (2018).

\bibitem{bib:30} S. A. Alkhaleefah, I. R. Kayumov and S. Ponnusamy,
“On the Bohr inequality with a fixed zero coefficient,” Proc.
Amer. Math. Soc. \textbf{147} (12), 5263--5274 (2019).

\bibitem{BhowDas-18} B. Bhowmik and N. Das,
“Bohr phenomenon for subordinating families of certain univalent functions,” J. Math. Anal. Appl. \textbf{462} (2), 1087--1098 (2018).

\bibitem{bib:21} R. M. Ali, R. W. Barnard, and A. Yu. Solynin,
“A note on the Bohr's phenomenon for power series,” J. Math.
Anal. Appl. \textbf{449} (1), 154--167 (2017).

\bibitem{bib:24} Y. Abu-Muhanna and R. M. Ali,
“Bohr's phenomenon for analytic functions and the hyperbolic metric,” Math. Nachr. \textbf{286} (11-12), 1059--1065 (2013).

\bibitem{bib:25} Y. Abu-Muhanna, R.M. Ali,  Z.C. Ng, and  S. F. M. Hasni,
“Bohr radius for subordinating families of analytic functions and
bounded harmonic mappings,” J. Math. Anal. Appl. \textbf{420}
(1), 124--136 (2014).

\bibitem{bib:26} Y. Abu-Muhanna and R. M. Ali,
“Bohr's phenomenon for analytic functions into the exterior of a
compact convex body,” J. Math. Anal. Appl. \textbf{379} (2),
512--517 (2011).

\bibitem{bib:31} M.-S. Liu, Y.-M. Shang and J.-F. Xu,
“Bohr-type inequalities of analytic functions,”
J. Inequal. Appl. \textbf{345}, 13 pp (2018).

\bibitem{bib:11} H. P. Boas and D. Khavinson,
“Bohr's power series theorem in several variables,” Proc. Am. Math. Soc. \textbf{125} (10), 2975--2979 (1997).

\bibitem{bib:12} L. Aizenberg,
“Multidimensional analogues of Bohr's theorem on power series,”
Proc. Amer. Math. Soc. \textbf{128}, 1147--1155 (2000).

\bibitem{bib:13} L. Aizenberg, “Generalization of results about the Bohr radius for power series,”
Stud. Math. \textbf{180}, 161--168 (2007).

\bibitem{bib:14} L. Aizenberg and N. Tarkhanov,
“A Bohr phenomenon for elliptic equations,” Proc. Amer. Math.
Soc. \textbf{82} (3), 385--401 (2001).

\bibitem{bib:16} I. R. Kayumov and S. Ponnusamy,
“Bohr--Rogosinski radius for analytic functions,”
Preprint. See {\tt https://arxiv.org/pdf/1708.05585.pdf}

\bibitem{AiElSh-05} %{bib:04}
L. Aizenberg, M. Elin, and D. Shoikhet, “On the Rogosinski radius for
holomorphic mappings and some of its applications,” Studia
Math. \textbf{168} (2), 147--158 (2005).

\bibitem{bib:15} I. R. Kayumov, S. Ponnusamy and N. Shakirov,
“Bohr radius for locally univalent harmonic mappings,” Math. Nachr.  \textbf{291}, 1757--1768 (2018).

\bibitem{bib:19} S. Evdoridis, S. Ponnusamy and A. Rasila,
“Improved Bohr's inequality for locally univalent harmonic mappings,”
Indag. Math. (N.S.)  \textbf{30}, 201--213 (2019).

\bibitem{bib:17} I. R. Kayumov and S. Ponnusamy,
“Bohr inequality for odd analytic functions,”
Comput. Methods Funct. Theory \textbf{17}, 679--688 (2017).

\bibitem{AAP2016} Y. Abu Muhanna, R. M. Ali, and S. Ponnusamy,
“On the Bohr inequality,” In “Progress in Approximation Theory
and Applicable Complex Analysis” (Edited by N.K. Govil et al. ),
Springer Optimization and Its Applications \textbf{117}, 265--295
(2016).

\bibitem{bib:32} L. Landau,
“Abschatzung der Koeffizientensumme einer Potenzreihe,”
Arch. Math. Physik \textbf{21}, 250--255 (1913).

\bibitem{bib:29} St. Ruscheweyh,
“Two remarks on bounded analytic functions,”
Serdica, Bulg. math. publ.  \textbf{11}, 200--202 (1985).

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

%\bibitem{bib:27} J. G. Clunie and T. Sheil-Small, Harmonic univalent functions,
%\emph{Ann. Acad. Sci. Fenn. Ser. A. I.} \textbf{9} (1984), 3--25.
%
%\bibitem{bib:28} P. Duren,
%Harmonic mappings in the plane,
%\emph{Cambridge University Press, Cambridge}, (2004), 226 p.

%I. R. Kayumov and S. Ponnusamy, Bohr’s inequality for analytic functions $\sum_k b_kz^{kp+m}$ and harmonic functions,
%(2017), Preprint, arXiv:1708.05578

%\bibitem{bib:23} R. Balasubramanian, B. Calado, H. Queff\'elec,
%The Bohr inequality for ordinary Dirichlet series,
%\emph{Stud. Math.} \textbf{175(3)} (2006), 285–-304.

%\bibitem{bib:07}
%S. Sidon, \"Uber einen Satz von Herrn Bohr,
%\emph{Math. Z.} \textbf{26(1)} (1927), 731-–732.

%\bibitem{bib:08} M. Tomi\'c, Sur un th\'eor\'eme de H. Bohr,
%\emph{Math. Scand.} \textbf{11} (1962), 103–-106.

%\bibitem{bib:10} V. I. Paulsen and D. Singh, A Simple Proof of Bohr’s Inequality,
%\emph{Conference Proceedings}, to appear.

\end{thebibliography}

\documentclass[
11pt,%
tightenlines,%
twoside,%
onecolumn,%
nofloats,%
nobibnotes,%
nofootinbib,%
superscriptaddress,%
noshowpacs,%
centertags]%
{revtex4}
\usepackage{ljmarxiv}
\usepackage{dsfont}
\usepackage{placeins}

\newtheorem{cor}{Corollary}
\newtheorem{lem}{Lemma}

\newtheorem{que}{Question}
\newtheorem{rem}{Remark}
\newtheorem{prob}{Problem}
\newtheorem{conj}{Conjecture}
\theoremstyle{definition}
\newtheorem{example}[equation]{Example}

\newcounter {own}
\def\theown {\thesection       .\arabic{own}}

\newenvironment{pf}[1][]{%
 \vskip 3mm
 \noindent
 \ifthenelse{\equal{#1}{}}%
  {{\slshape Proof. }}%
  {{\slshape #1.} }%
 }%
{\qed\bigskip}

\newcounter{alphabet}

\newenvironment{Thm}[1][]{\refstepcounter{alphabet}%
\bigskip%
\noindent%
{\bf Theorem \Alph{alphabet}}%
\ifthenelse{\equal{#1}{}}{}{ (#1)}%
{\bf .} \itshape}{\vskip 8pt}

\newcommand{\IN}{{\mathbb N}}
\newcommand{\IC}{{\mathbb C}}
\newcommand{\ID}{{\mathbb D}}

\def\be{\begin{equation}}
\def\ee{\end{equation}}

\newcommand{\bee}{\begin{enumerate}}
\newcommand{\eee}{\end{enumerate}}

\newcommand{\blem}{\begin{lem}}
\newcommand{\elem}{\end{lem}}
\newcommand{\bthm}{\begin{theorem}}
\newcommand{\ethm}{\end{theorem}}
\newcommand{\bcor}{\begin{cor}}
\newcommand{\ecor}{\end{cor}}
\newcommand{\beg}{\begin{example}}
\newcommand{\eeg}{\end{example}}
\newcommand{\bques}{\begin{que}}
\newcommand{\eques}{\end{que}}
\newcommand{\begs}{\begin{examples}}
\newcommand{\eegs}{\end{examples}}
\newcommand{\bdefe}{\begin{defin}}
\newcommand{\edefe}{\end{defin}}
\newcommand{\bprob}{\begin{prob}}
\newcommand{\eprob}{\end{prob}}
\newcommand{\bei}{\begin{itemize}}
\newcommand{\eei}{\end{itemize}}

\newcommand{\bcon}{\begin{conj}}
\newcommand{\econ}{\end{conj}}
\newcommand{\bcons}{\begin{conjs}}
\newcommand{\econs}{\end{conjs}}
\newcommand{\bprop}{\begin{propo}}
\newcommand{\eprop}{\end{propo}}
\newcommand{\br}{\begin{rem}}
\newcommand{\er}{\end{rem}}
\newcommand{\brs}{\begin{rems}}
\newcommand{\ers}{\end{rems}}
\newcommand{\bo}{\begin{obser}}
\newcommand{\eo}{\end{obser}}
\newcommand{\bos}{\begin{obsers}}
\newcommand{\eos}{\end{obsers}}
\newcommand{\bpf}{\begin{pf}}
\newcommand{\epf}{\end{pf}}
\newcommand{\ba}{\begin{array}}
\newcommand{\ea}{\end{array}}
\newcommand{\beq}{\begin{eqnarray}}
\newcommand{\beqq}{\begin{eqnarray*}}
\newcommand{\eeq}{\end{eqnarray}}
\newcommand{\eeqq}{\end{eqnarray*}}

\newcommand{\ra}{\rightarrow}

\newcommand{\ds}{\displaystyle}

\begin{document}

\titlerunning{Bohr--Rogosinski inequalities} 
\authorrunning{Seraj A.~Alkhaleefah et al.} 

\title{Bohr--Rogosinski inequalities
for bounded analytic functions}

\author{\firstname{Seraj A.}~\surname{Alkhaleefah}}
\email[E-mail: ]{s.alkhaleefah@gmail.com}
\affiliation{N.I. Lobachevskii Institute of Mathematics and Mechanics, Kazan (Volga Region) Federal University, Kremlevskaya ul. 18, Kazan, Tatarstan, 420008 Russia}

\author{\firstname{Ilgiz R.}~\surname{Kayumov}}
\email[E-mail: ]{ikayumov@gmail.com}
\affiliation{N.I. Lobachevskii Institute of Mathematics and Mechanics, Kazan (Volga Region) Federal University, Kremlevskaya ul. 18, Kazan, Tatarstan, 420008 Russia}

\author{\firstname{Saminathan}~\surname{Ponnusamy}}
\email[E-mail: ]{samy@iitm.ac.in}
\affiliation{Department of Mathematics,
Indian Institute of Technology Madras,
Chennai-600 036, India}
%


\subclass{Primary: 30A10, 30B10; 30C62, 30H05, 31A05, 41A58; Secondary:  30C75, 40A30} 

\keywords{Bounded analytic function, Bohr inequality, Bohr radius, Rogosinski inequality, Rogosinski radius, harmonic mappings.} 

\begin{abstract}
In this paper we first consider another version of the Rogosinski inequality for analytic functions
$f(z)=\sum_{n=0}^\infty a_nz^n$ in the unit disk $|z| < 1$, in which we replace the coefficients $a_n$ $(n= 0,1,\ldots ,N)$ of
the power series by the derivatives $f^{(n)}(z)/n!$ $(n= 0,1,\ldots ,N)$. Secondly, we
obtain improved versions of the classical Bohr inequality and Bohr's inequality for the harmonic mappings
of the form $f = h + \overline{g}$, where the analytic part $h$ is bounded by $1$ and
that $|g'(z)| \le k|h'(z)|$ in $|z| < 1$ and for some $k \in [0,1]$.
\end{abstract}

\maketitle


\section{Introduction}
Let $\ID=\{z\in \IC:\, |z| < 1 \}$  denote the open unit disk, and
${\mathcal A}$ denote the space of analytic functions in $\ID$ with
the topology of uniform convergence on compact sets. Define
${\mathcal B}=\{f\in {\mathcal A}:\, \mbox{$|f(z)|<1$ in $\ID$}\}$.
Then the Bohr radius is the largest number $r>0$ such that if $f\in
{\mathcal B} $ has the power series expansion
$f(z)=\sum_{n=0}^\infty a_n z^n$, then
$\sum_{n=0}^{\infty}|a_n|\,|z|^n \leq 1 ~\mbox{ for }~ |z|\leq r
$
which is called the classical Bohr inequality for the family
${\mathcal B}$. Rogosinski radius is the largest number $r>0$ such
that, under the previous assumptions, $\left|S_{N}(z) \right|<1
~\mbox{ for }~ |z|<r,
$
where $S_{N}(z)=\sum_{n=0}^{N}a_nz^n$ $(N\geq 0)$ denote the partial sums of $f$. This inequality is called
the classical Rogosinski inequality for the family ${\mathcal B}$.

 If $\textbf{B}$ and $\textbf{R}$ denote the Bohr radius and the
Rogosinski radius, respectively, then because $\left |S_{N}(z)\right |\leq \sum_{n=0}^{N}|a_n|\,|z|^n\leq \sum_{n=0}^{\infty}|a_n|\,|z|^n,$
it is clear that $\textbf{B}\leq \textbf{R}$. In fact the following two classical results are well-known.

\begin{Thm} Suppose that $f\in {\mathcal B} $. Then we have
{\rm $\textbf{B}=1/3$, and  {\rm (see Rogosinski \cite{Rogo-9} and also \cite{LanDai-86,SchSz-25})} $\textbf{R}=1/2$.}
\end{Thm}



There is a long history about the consequences of Bohr's inequality, in particular.
Indeed, Bohr \cite{Bohr-14} discovered that $\textbf{B}\geq 1/6$ and
the fact that $\textbf{B}=1/3$ was obtained independently by M. Riesz, I. Schur and N. Weiner.
Extensions and modifications of Bohr's result can be found from
\cite{bib:22,PPS2002,bib:09} and the recent articles
\cite{bib:18,bib:30,BhowDas-18,
bib:21,bib:24,bib:25,bib:26,bib:31}). We refer to
\cite{bib:11,bib:12,bib:13,bib:14} for the extension of the Bohr
inequality to several complex variables. More recently, Kayumov and
Ponnusamy \cite{bib:16} introduced and investigated
Bohr--Rogosinski's radii for the family $\mathcal B$, and they
discussed Bohr--Rogosinski's radius for the class of subordinations.
In \cite{AiElSh-05}, Aizenberg, et al. generalized the Rogosinski
radius for holomorphic mappings of the open unit polydisk into an
arbitrary convex domain. In \cite{bib:15},  Kayumov et al.
investigated Bohr's radius for complex-valued harmonic mappings that
are locally univalent in $\ID$. Several improved versions of Bohr's
inequality were given by Kayumov and Ponnusamy in \cite{bib:16} and
these were subsequently followed by  Evdoridis et al.  \cite{bib:19}
to obtain improved versions of Bohr's inequality for the class of
harmonic mappings. In \cite{bib:17}, Kayumov and Ponnusamy discussed
Bohr's radius for the class of analytic functions $g$, when $g$ is
subordinate to a member of the class of odd univalent functions. For
more information about Bohr's inequality and further  related works,
we refer the reader to the recent survey article \cite{AAP2016} and
the references therein.

In this paper we shall introduce and investigate another version of
the Rogosinski inequality for  analytic functions defined on the
unit disk $\mathbb{D}$ by substituting the derivatives of the
analytic function instead of the coefficients of its power series.
We shall also introduce and study several new versions of the
classical Bohr's inequality.

\section{An improved version of the classical Rogosinski inequality}
What could happened to the partial sums of the analytic function in
the unit  disk if we replaced the coefficients
$a_0,a_1,\ldots,a_{N-1}$ by the functions
$f(z),f'(z),\ldots,f^{(N-1)}(z)$? In this section we give an answer
in the following form.

\bthm\label{AKS2-th1}
  Suppose that $f\in {\mathcal B}$ and $f(z)= \sum_{n=0}^\infty a_n z^n$. Then
\begin{equation*}
\left|\sum_{k=0}^n\frac{f^{(k)}(z)}{k!}z^k\right|\le \sum_{k=0}^n \binom{-\frac12}{k}^2 ~\mbox{ for all $\ds |z|\leq r\le\frac12$}.
\end{equation*}
\ethm
\bpf
To prove this theorem we will use a modification of Landau's method (see \cite{bib:32} and \cite[p.~26]{LanDai-86}).

We consider the function $g:\, \ID \to \ID$ defined by $g(\zeta)=f(\alpha (\zeta +1))$, where $|\alpha|\leq 1/2$, and use the
substitution $\xi= D(\zeta) = \alpha (\zeta +1)$. In view of the Cauchy integral formula, integration
along a circle $\gamma$ around the origin lying in its neighborhood, we have
\begin{equation*}
\frac{f^{(k)}(\alpha)}{k!}  = \frac1{2\pi i}\int_{D(\gamma)} \frac{f(\xi)}{(\xi -\alpha)^{k+1}}\,d\xi
= \frac1{2\pi i \alpha^k}\int_{\gamma } \frac{g(\zeta)}{\zeta^{k+1}}\,d\zeta
\end{equation*}
and thus,  we can write
\begin{equation}\label{inqty_3}
\sum_{k=0}^n\frac{f^{(k)}(\alpha)}{k!}\alpha^k = \frac1{2\pi i}\int_{\gamma} g(\zeta) \left (\sum_{k=0}^n \frac{1}{\zeta^{k+1}}\right )
 d\zeta = \frac1{2\pi i}\int_{\gamma} \frac{g(\zeta)}{\zeta^{n+1}} \left (\sum_{k=0}^n \zeta^k\right ) d\zeta.
\end{equation}
Set
$1+\zeta+\zeta^2+\zeta^3+\dots = (1-\zeta)^{-1}  =  K^2(\zeta)= (K_n(\zeta))^2 +O(\zeta ^{n+1}),
$
where we write
$$
K(\zeta) =  (1-\zeta)^{-1/2}  =\sum_{k=0}^\infty \binom{-\frac12}{k} (-\zeta)^k ~\mbox{ and }~
K_n(\zeta) = \sum_{k=0}^n \binom{-\frac12}{k} (-\zeta)^k.
$$

In view of the above observations, \eqref{inqty_3} reduces to
\begin{equation}\label{eqty_1}
\sum_{k=0}^n\frac{f^{(k)}(\alpha)}{k!}\alpha^k = \frac1{2\pi i}\int_{\gamma} \frac{g(\zeta)}{\zeta^{n+1}} (K_n(\zeta))^2 \,d\zeta
\end{equation}
and therefore, with $\zeta=|\zeta|e^{i\phi}$, and $|g(\zeta)|\le1$ for all $|\alpha|\le1/2$ and $|\zeta|\le1$, we have
\begin{eqnarray*}
\left|\sum_{k=0}^n\frac{f^{(k)}(\alpha)}{k!}\alpha^k\right|
&\le& \frac1{2\pi}\int_0^{2\pi} \frac{1}{|\zeta|^{n+1}} |K_n(\zeta)|^2|\zeta| \,d\phi
\, =\,  \frac1{|\zeta|^n} \sum_{k=0}^n \binom{-\frac12}{k}^2 |\zeta|^{2k}.
\end{eqnarray*}
Allowing $|\zeta|\ra 1$, we get
\begin{equation*}
\left|\sum_{k=0}^n\frac{f^{(k)}(\alpha)}{k!}\alpha^k\right|
\le   \sum_{k=0}^n \binom{-\frac12}{k}^2 ~\mbox{ for all $ \ds |\alpha|\le \frac12$}
\end{equation*}
which completes the proof of the theorem.
\epf

\section{Improved versions of the classical Bohr's inequality}

For $f\in {\mathcal B}$ and $f(z)= \sum_{n=0}^\infty a_n z^n$, the following inequalities due to Schwarz-Pick will be used frequently: for $|z|=r<1$,
\begin{equation}\label{AKP-SchPick}
|f(z)|\le \frac{r+a}{1+ar} ~\mbox{ and }~ |f'(z)|\le \frac{1-|f(z)|^2}{1-|z|^2},
\end{equation}
where $|a_0|=a\in [0,1)$. Also, it is well-known that the Taylor coefficients of  $f\in {\mathcal B}$ satisfy the inequalities:
\begin{equation}\label{AKP-SchPick2}
|a_k|\le 1-a^2 ~\mbox{ for each $k\ge 1$.}
\end{equation}
More generally, we have (\cite{bib:29}) the sharp estimate
\be\label{AKPs-Rusheq2}
\frac{|f^{(k)}(z)|}{k!}\le \frac{1-|f(z)|^2}{(1-|z|)^k(1+|z|)} ~\mbox{ for each $k\ge 1$ and $z\in\ID$,}
\ee
which in particular gives second inequality in \eqref{AKP-SchPick}, and \eqref{AKP-SchPick2} by setting
$z=0$ in \eqref{AKPs-Rusheq2}. In the following we also assume that $m\in \IN$, and the idea of
replacing $a_k$ by $\frac{f^{(k)}(z)}{k!}$ is used in \cite{bib:31}. But our concern here is slightly different
from theirs.

\bthm\label{AKP2-th1}
Suppose that $f\in {\mathcal B}$ and $f(z)= \sum_{n=0}^\infty a_n z^n$. Then
\begin{equation}\label{AKP-eq1}
A_f(z):=|f(z^m)|+|z^m|\,|f'(z^m)| +\sum_{k=2}^\infty |a_k|r^k \le 1~\mbox{ for all }~ r \le R_{m,1},
\end{equation}
where $R_{m,1}$ is the maximal positive root of the equation $\varphi_m(r)=0$ with
\begin{equation}\label{AKP-eq2}
\varphi_m(r)=(1-r)(r^{2m}+2r^m-1)+2r^2(1+r^m)^2
\end{equation}
and the constant $R_{m,1}$ cannot be improved.
\ethm
\begin{table}[!htbp]
\begin{center}
\begin{tabular}{|l|l|}
\hline
 {\bf $m$}   & $R_{m,1}$ \\
  \hline
1  & 0.280776   \\
  \hline
2  & 0.39149   \\
  \hline
3  & 0.441112   \\
  \hline
4  & 0.467644   \\
  \hline
5  & 0.482442   \\
  \hline
\end{tabular}
\end{center}
\caption{$R_{m,1}$ is the maximal positive root of the equation
$(1-r) (r^{2m}+2r^m-1) + 2r^2(1+r^m)^2=0$\label{AKP2-th1-T1}}
\end{table}
\bpf
Let $f\in {\mathcal B}$ and $|a_0|=a\in [0,1)$. 
 It is a simple exercise to see that for $0\leq x\leq x_0 ~(\leq 1)$ and $0\leq \alpha \leq 1/2$, we have
$b(x):=x+\alpha (1-x^2)\leq b(x_0). $ This simple fact will be used
in the later theorems also. Using this inequality and
\eqref{AKP-SchPick2}, we easily obtain from \eqref{AKP-SchPick} and
\eqref{AKP-eq1}  that
\begin{align*}
A_f(z) &\le |f(z^m)| + \frac{r^m}{1-r^{2m}}(1-|f(z^m)|^2)
+(1-a^2)\frac{r^2}{1-r}\\
& \le \frac{r^m+a}{1+ar^m} +
\frac{r^m}{1-r^{2m}}\left[1-\left(\frac{r^m+a}{1+ar^m}\right)^2\right]
+(1-a^2)\frac{r^2}{1-r}\\
&= \frac{r^m+a}{1+ar^m} + (1-a^2)\frac{r^m}{(1+ar^m)^2} +(1-a^2)\frac{r^2}{1-r}\\
&= 1-\frac{(1-a)(1-r^m)}{1+ar^m} + (1-a^2)\frac{r^m}{(1+ar^m)^2}
+(1-a^2)\frac{r^2}{1-r}
=1+\frac{(1-a)\varphi_m(a,r)}{(1+ar^m)^2(1-r)},
\end{align*}
where
\begin{align*}
\varphi_m(a,r) &= -(1-r^m)(1+ar^m)(1-r) +(1+a)r^m(1-r) +
r^2(1+a)(1+ar^m)^2 \\
&= (1-r)(ar^{2m}+2r^m-1) + r^2(1+a)(1+ar^m)^2.
\end{align*}
The second inequality above is justified because of the fact that
$\frac{r^m}{1-r^{2m}}\leq \frac{1}{2}$ for $r \leq
\sqrt[m]{\sqrt2-1}.$ Also,  $R_{m,1}\leq \sqrt[m]{\sqrt2-1}$, where
$R_{m,1}$ is as in the statement. Now, since $\varphi_m(a,r) $ is an
increasing function of $a$ in $[0,1)$, it follows that
$$\varphi_m(a,r)\leq \varphi_m(1,r)=(1-r)(r^{2m}+2r^m-1) + 2r^2(1+r^m)^2= \varphi_m(r),
$$
where $\varphi_m(r)$ is given by (\ref{AKP-eq2}). Clearly,
$A_f(z)\leq 1$ if $\varphi_m(r)\le 0$, which holds for $r\le
R_{m,1}$.

To show the sharpness of the radius $R_{m,1}$, we let $a \in[0,1)$ and consider the function
\begin{equation}\label{AKP2-eqex1}
f(z)=\frac{a+z}{1+az}=a+(1-a^2)\sum_{k=1}^\infty (-a)^{k-1} z^k, \ \ z\in \ID
\end{equation}
so that
$$\frac{f^{(k)}(z)}{k!}=(1-a^2)\frac{ (-a)^{k-1}}{(1+az)^{k+1}} ~\mbox{ for  $k\ge 1$ and $z\in\ID$.}
$$
For this function, we observe that for $z=r$ and $a\in [0,1)$,
\beq\label{AKP-eq3}
|f(z^m)|+|z^m|\,|f'(z^m)| +\sum_{k=2}^\infty |a_k|r^k
&=& \frac{a+2r^m+ar^{2m}}{(1+ar^m)^2}+(1-a^2)\frac{ar^2}{1-ar}\nonumber\\
&=& 1+\frac{(1-a)P_m(a,r) }{(1+ar^m)^2(1-ar)}, \eeq where
$P_m(a,r)=(1-ar)(ar^{2m}+2r^m-1) + ar^2(1+a)(1+ar^m)^2
$
and the last expression \eqref{AKP-eq3} is larger than $1$ if and
only if $P_m(a,r)>0$. By a simple calculation, we find that
$$\frac{\partial P_m(a,r)}{\partial a}= r^{2m} +2a(r^2-r^{2m+1}) +r(1+r-2r^m)+3a^2r^{2m+2}+4ar^{m+2}+2ar^2+4a^3r^{2m+2}+6a^2r^{m+2}
$$
which is clearly non-negative for each $r\in [0,1)$ and thus, for each $r\in [0,1)$, $P_m(a,r)$ is an increasing function of $a$. This
fact gives
$$0< r^{2m} +r(1+r-2r^m) =P_m(0,r)\leq P_m(a,r)\le P_m(1,r)= \varphi_m(r),
$$
where $\varphi_m(r)$ is given by (\ref{AKP-eq2}).
Therefore, the right hand side of (\ref{AKP-eq3}) is smaller than or equal to $1$ for all $a\in [0,1)$, only in the case $r\le R_{m,1}$.
Finally, it also suggests that the right hand side of (\ref{AKP-eq3}) is larger than $1$ if $r > R_{m,1}$. This completes the proof.
\epf

\br
In Tables \ref{AKP2-th1-T1} and \ref{AKP2-th2-T2}, we listed the values of $R_{m,1}$ and $R_{m,2}$ for certain values of $m$.
If we allow $m\to\infty$ in Theorem \ref{AKP2-th1} (resp. Theorem \ref{AKP2-th2} below),
we note that $R_{m,1}\to 1/2$ (resp. $R_{m,2}\to 1/2$ below) and thus if $f\in {\mathcal B}$, then we have the
inequality
$$
|f(0)|+\sum_{k=2}^\infty \left|\frac{f^{(k)}(0)}{k!}\right|r^k \le 1 ~\mbox{ for all $ r\le 1/2$,}
$$
and the number $1/2$ is sharp.
\er

\bthm\label{AKP2-th2}
If $f\in {\mathcal B}$, then
\begin{equation}\label{AKP-eq5}
B_f(z):= |f(z^m)| + \sum_{k=2}^\infty \left|\frac{f^{(k)}(z^m)}{k!}\right|r^k \le 1~\mbox{for all $r \le R_{m,2}$},
\end{equation}
where $R_{m,2}$ is the minimum positive root of the equation $\psi _m(r)=0$ with
\begin{equation}\label{AKP-eq6}
\psi _m(r)=2r^2-(1-r^{2m})(1-r^m-r)
\end{equation}
and the constant $R_{m,2}$ cannot be improved.
\ethm
\begin{table}[!htbp]
\begin{center}
\begin{tabular}{|l|l|}
\hline
 {\bf $m$}   & $R_{m,2}$ \\
  \hline
1  & 0.355416   \\
  \hline
2  & 0.430586   \\
  \hline
3  & 0.464327 \\
  \hline
4  &  0.481418 \\
  \hline
5  & 0.490359  \\
  \hline
\end{tabular}
\end{center}
\caption{$R_{m,2}$ is the maximal positive root of the equation
$2r^2-(1-r^{2m})(1-r^m-r)=0$\label{AKP2-th2-T2}}
\end{table}
\bpf
As before we let $f\in {\mathcal B}$ and $a=|a_0|$.
By assumption, \eqref{AKP-SchPick} and \eqref{AKPs-Rusheq2} (with $z^m$ in place of $z$), we have
\begin{eqnarray*}
B_f(z)&\le& |f(z^m)|+ \frac{1-|f(z^m)|^2}{1+r^m}\sum_{k=2}^\infty \left (\frac{r}{1-r^m}\right )^k \\
&= & |f(z^m)|+ \frac{r^2}{(1-r^{2m})(1-r^m-r)}(1-|f(z^m)|^2)\\
&\le& \frac{r^m+a}{1+ar^m} + \frac{r^2}{(1-r^{2m})(1-r^m-r)}\left[1-\left(\frac{r^m+a}{1+ar^m}\right)^2\right] \\
&=& 1-\frac{(1-a)(1-r^m)}{1+ar^m} +
\frac{r^2(1-a^2)}{(1+ar^m)^2(1-r^m-r)}=
1+\frac{(1-a)\psi_m(a,r)}{(1+ar^m)^2(1-r^m-r)},
\end{eqnarray*}
where
\beqq
\psi_m(a,r) &=& -(1-r^m)(1+ar^m)(1-r^m-r)+r^2(1+a).
\eeqq
The second inequality above is a consequence of our earlier observation used in Theorem \ref{AKP2-th1} but this time with
$\alpha = r^2/[(1-r^{2m})(1-r^m-r)]$. It is a simple exercise to see that $\psi_m(a,r)$, for each $m\geq 1$,
is an increasing function of $a$ in $[0,1)$, and thus, it follows that
$
\psi_m(a,r) \leq  \psi_m(1,r) =\psi_m(r),
$
where $\psi_m(r)$ is given by \eqref{AKP-eq6}.
Clearly, $\psi_m(a,r)\le0$ if $\psi_m(r)\le 0$, which holds for $r\le R_{m,2}$, where
$R_{m,2}$ is the minimum positive root of the equation $\psi_m(r)= 0$. Thus, $B_f(z)\leq 1$
for $r\le R_{m,2}$ and the inequality \eqref{AKP-eq5} follows.

To show the sharpness of the radius $R_{m,2}$, we let $a \in(0,1)$ and consider the function
$g(z)=f(-z)$, where $f$ is given by \eqref{AKP2-eqex1}. Then for $g$, we easily have
$$\frac{g^{(k)}(z)}{k!} =-(1-|a|^2)\frac{a^{k-1}}{(1-az)^{k+1}},  \ \ z\in \ID.
$$
Now, we choose $a$ as close to $1$ as we please and set $z=r<\sqrt[m]a$. By a simple calculation,
the corresponding $B_g(z)$  takes the form
\beq\label{AKP-eq7}
\hspace{-.8cm}B_g(z) &=& \frac{a-r^m}{1-ar^m} + \frac{ar^2(1-a^2)}{(1-ar^m)^2(1-ar^m-ar)}\nonumber\\
&=& 1-\frac{(1-a)(1+r^m)}{1-ar^m} +
\frac{ar^2(1-a^2)}{(1-ar^m)^2(1-ar^m-ar)}=
1+\frac{(1-a)P_m(a,r)}{(1-ar^m)^2(1-ar^m-ar)},
\eeq
where
$$P_m(a,r)=a(1+a)r^2-(1+r^m)(1-ar^m)(1-ar^m-ar).
$$
Clearly, $B_g(z)<1$ if and only if $P_m(a,r)<0$. By elementary calculations, we find that
\beqq
\frac{\partial P_m(a,r)}{\partial a}&=& 2a[r^2-r^m(1+r^{m})(r^{m}+r)]+r^2+(1+r^{m})(2r^m+r)\\
&=&2a(r^2-r^{2m}-r^{m+1}-r^{3m}-r^{2m+1})+(2r^{2m}+r^{m+1}+2r^m+r^2+r)
\eeqq
which is easily seen to be greater than or equal to $0$ for any $r\in [0,1)$ and $m\ge 1$. Consequently,
$$P_m(a,r)\le P_m(1,r)=\psi _m(r) =2r^2-(1-r^{2m})(1-r^m-r).
$$
Therefore, the expression on the right of (\ref{AKP-eq7}) is smaller than or equal to $1$ for all $a\in (0,1)$, only in the case when $r\le R_{m,2}$.
Finally, it also suggests that $a\to 1$ in the right hand side of (\ref{AKP-eq7}) shows that  the expression (\ref{AKP-eq7}) is larger than
$1$ if and only if $r > R_{m,2}$. This completes the proof.
\epf

\bthm\label{AKP2-th3}
Suppose that $f\in {\mathcal B}$ and $f(z)= \sum_{n=0}^\infty a_n z^n$. Then
\begin{equation}\label{AKP-eq15}
C_f(z):=|f(z^m)|+|z|\,|f'(z^m)| +\sum_{k=2}^\infty |a_k|r^k \le 1\quad\text{for all}\quad r \le R_{m,3},
\end{equation}
where $R_{m,3}$ is the maximal positive root of the equation $\Phi_m (r)=0$ with
\begin{equation}\label{AKP-eq16}
\Phi_m (r)= 3r-1 + r^m\left[2r^2(r^m+2)+r^m(1-r)\right]
\end{equation}
and the constant $R_{m,3}$ cannot be improved.
\ethm
\begin{table}[!htbp]
\begin{center}
\begin{tabular}{|l|l|}
\hline
 {\bf $m$}   & $R_{m,3}$ \\
  \hline
1  & 0.280776   \\
  \hline
2  & 0.316912   \\
  \hline
3  & 0.327911 \\
  \hline
4  & 0.33152 \\
  \hline
5  & 0.332726  \\
  \hline
\end{tabular}
\end{center}
\caption{$R_{m,3}$ is the maximal positive root of the equation
$3r-1 + r^m[2r^2(r^m+2)+r^m(1-r)]=0$ \label{AKP2-th3-T3}}
\end{table}
\bpf
As in the proofs of Theorems \ref{AKP2-th1} and \ref{AKP2-th2}, it follows from \eqref{AKP-SchPick}, \eqref{AKP-SchPick2} and \eqref{AKPs-Rusheq2} that
\begin{eqnarray*}
C_f(z) &\le& |f(z^m)| + \frac{r}{1-r^{2m}}(1-|f(z^m)|^2) +(1-a^2)\frac{r^2}{1-r} \\
&\le& \frac{r^m+a}{1+ar^m} + \frac{r}{1-r^{2m}}\left[1-\left(\frac{r^m+a}{1+ar^m}\right)^2\right] +(1-a^2)\frac{r^2}{1-r} \\
&=& 1-\frac{(1-a)(1-r^m)}{1+ar^m} + (1-a^2)\frac{r}{(1+ar^m)^2} +(1-a^2)\frac{r^2}{1-r} \\
&=& \frac{a+2r^m+ar^{2m}}{(1+ar^m)^2}+(1-a^2)\frac{r^2}{1-r}=1+\frac{(1-a)\Phi_m(a,r)}{(1+ar^m)^2(1-r)}
\end{eqnarray*}
where
$$
\Phi_m(a,r) =
-(1-r^m)(1+ar^m)(1-r) + r(1+a)(1-r) +  r^2(1+a)(1+ar^m)^2
$$
$$
= r(1+a) + ar^{m+2}(1+a)(2+ar^m)-(1-r^m)(1+ar^m)(1-r)
$$
$$
\le \Phi_m(1,r) =\Phi_m(r) 
$$
because $\Phi_m(a,r) $ is seen to be an increasing function of $a$ in
$[0,1)$, and $\Phi_m(r)$ is given by \eqref{AKP-eq16}. Note that the
second inequality above holds since $\max \frac{2r}{1-r^{2m}}<1$ and
so for any $r <R_m$, where $R_m$ is the maximal positive root of the
equation $2r-(1-r^{2m})=0$, and $R_{m,3}<R_m$ for $m\in \IN$, where
$R_{m,3}$ is the maximal positive root of the equation $\Phi_m
(r)=0$. Since $\Phi_m(r)\le0$  for $r\le R_{m,3}$, we obtain
$C_f(z)\leq 1$ for $r\le R_{m,3}$ and the desired inequality
\eqref{AKP-eq15} follows.

It remains to show the sharpness of the radius $R_{m,3}$. To do this we let $a \in[0,1)$ and consider the function
$f$ is given by \eqref{AKP2-eqex1}.
For this function, we observe that for $z=r$,
\begin{equation}\label{AKP-eq17}
C_f(z) = \frac{(r^m+a)(1+ar^m)+r(1-a^2)}{(1+ar^m)^2}+(1-a^2)\frac{ar^2}{1-ar}= 1+\frac{(1-a)Q_m(a,r)}{(1+ar^m)^2(1-r)},
\end{equation}
where
$$Q_m(a,r)=r(1+a) + a^2r^{m+2}(1+a)(2+ar^m)-(1-r^m)(1+ar^m)(1-ar).
$$
We see that   $C_f(z)>1$ for $a \in[0,1)$ if and only if $Q_m(a,r) >0$.
By a simple calculation, we find that $Q_m(a,r)$ is an increasing function of $a$ in $[0,1)$
and therefore, we have
\begin{eqnarray*}
Q_m(a,r)\le Q_m(1,r) = 2r + 2r^{m+2}(2+r^m)-(1-r^m)(1+r^m)(1-r)=\Phi_m(r),
\end{eqnarray*}
where $\Phi_m(r)$ is given by \eqref{AKP-eq16}.
Therefore, the expression (\ref{AKP-eq17}) is smaller than or equal to $1$ for all $a\in [0,1)$, only  when
$r\le R_{m,3}$. Finally, it also suggests that $a\to 1$ in (\ref{AKP-eq17}) shows that the expression (\ref{AKP-eq17}) is
larger than $1$ if $r > R_{m,3}$. This completes the proof.
\epf

\br
In Table \ref{AKP2-th3-T3}, we listed the values of $R_{m,3}$ for certain values of $m$.
If we allow $m\to\infty$ in Theorem \ref{AKP2-th3}, we see that $R_{m,3}\to \frac13$ and hence we have the classical Bohr inequality
for $f\in {\mathcal B}$:
$$ |f(0)| +\sum_{k=1}^\infty |a_k|\,|z|^k \le 1 ~\mbox{ for all $r\le1/3$},
$$
and $1/3$ is sharp.
\er


\section{Two Improved versions of Bohr's inequality for harmonic mappings}

\bthm\label{AKP2-th4}
Suppose that $f(z)= h(z) + \overline{g(z)} =\sum_{n = 0}^\infty a_nz^n
+ \overline{\sum_{n = 1}^\infty b_nz^n}$ is a harmonic mapping of $\ID$ such that
$|g'(z)|\leq k|h'(z)|$ for some $k\in [0,1]$ and $h\in {\mathcal B}$. Then we have
\begin{equation}\label{AKP-eq9}
D_f(z):=
|h(z^m)|+\sum_{n=1}^\infty |a_n|r^n + \sum_{n=1}^\infty |b_n|r^n \le 1~\mbox{ for all $r \le R^k_{m,1}$},
\end{equation}
where $R^k_{m,1}$ is the maximal positive root of the equation $\lambda_m (r)=0$ with
\begin{equation}\label{AKP-eq10}
\lambda_m (r)= 2r(1+k)(1+r^m)-(1-r)(1-r^m)
\end{equation}
and the constant $R^k_{m,1}$ cannot be improved.
\ethm
\begin{table}[!htbp]
\begin{center}
\begin{tabular}{|l|l|}
\hline
 {\bf $m$}   & $R^1_{m,1}$ \\
  \hline
1  &  0.154701  \\
  \hline
2  & 0.188829   \\
  \hline
3  & 0.197544 \\
  \hline
4  & 0.199494 \\
  \hline
5  & 0.199898  \\
  \hline
\end{tabular}
\end{center}
\caption{$R^1_{m,1}$ is the maximal positive root of the equation
$4r(1+r^m)-(1-r)(1-r^m)=0$ \label{AKP2-th4-T4}}
\end{table}
\bpf
Recall that, as $h\in {\mathcal B}$ and $h(z)=\sum_{n = 0}^\infty a_nz^n$,
$$h(z)\prec  \frac{a_0+z}{1+\overline{a_0}z}=a_0+(1-|a_0|^2)\sum_{n=1}^\infty (-1)^{n-1}(\overline{a_0})^{n-1}z^n, \ \ z\in \ID,
$$
which gives \cite{BhowDas-18,bib:30}
\begin{eqnarray*}
\sum_{n=1}^{\infty}|a_{n}|r^{n}\leq (1-|a_0|^2)\sum_{n=1}^{\infty} |a_0|^{n-1}r^{n}= \frac{(1-a^2)r}{1-ar}~\mbox{ for all }\, r \leq\frac{1}{3},
\end{eqnarray*}
where $a=|a_0|\in[0,1)$. Moreover, by assumption, we obtain that $g'(z)\prec_q k h'(z)$ which quickly gives from  \cite{bib:30} that
\begin{eqnarray*}
\sum_{n=1}^{\infty}n|b_{n}|r^{n-1}\leq \sum_{n=1}^{\infty}k n\, |a_{n}|r^{n-1}~\mbox{ for all }\, r \leq\frac{1}{3}
\end{eqnarray*}
and integrating this with respect to $r$ gives
$$\sum_{n=1}^{\infty}|b_{n}|r^{n}\leq k\sum_{n=1}^{\infty}|a_{n}|r^{n}~\mbox{ for all }\, r \leq\frac{1}{3}.
$$
Here $\prec _q$ denotes the quasi-subordination. Using these and the first inequality in \eqref{AKP-SchPick} for $h(z)$ one can
obtain that for $|z|=r\le 1/3$,
\begin{align*}
D_f(z) &\le  \frac{r^m+a}{1+ar^m} + (1+k)r\frac{1-a^2}{1-ar}
=1+\frac{(1-a)\lambda_m(a,r)}{(1+ar^m)(1-ar)},
\end{align*}
where
$$\lambda_m(a,r) = r(1+k)(1+a)(1+ar^m)-(1-r^m)(1-ar),
$$
which is indeed an  increasing function of $a\in [0,1)$ so that
$\lambda_m(a,r) \leq \lambda_m(1,r)= \lambda_m(r),$ where $\lambda_m (r)$
is given by \eqref{AKP-eq10}. We see that $D_f(z)\leq 1$ if $\lambda
_m(r)\le0$, which holds for $r\le R^k_{m,1}$, where $R^k_{m,1}$ is
the maximal positive root of the equation  $\lambda_m (r)=0$. This
proves the inequality \eqref{AKP-eq9}.

Finally, to show the sharpness of the radius $R^k_{m,1}$,  we consider the function
\be\label{AKP-eq18}
f(z)=h(z) +\lambda k \overline{h(z)}, \quad h(z)= \frac{z+a}{1+az},
\ee
where  $\lambda \in (0,1]$. For this function, we get that (for $z=r$ and $a\in [0,1)$)
$$ D_f(z)  = \frac{r^m+a}{1+ar^m} + (\lambda k +1) r\frac{1-a^2}{1-ra}
$$
and the last expression shows the sharpness of $R^k_{m,1}$ with $\lambda \to 1$. This completes the proof of the theorem.
\epf

\br In Table \ref{AKP2-th4-T4}, we listed the values of $R^k_{m,1}$ for $k=1$ and for certain values of $m$.
When $m\to\infty$, we have from Theorem \ref{AKP2-th4} that $R^k_{m,1}\to \frac{2}{4k+6}$. Thus, under the hypotheses
of Theorem \ref{AKP2-th4}, we have
$$|h(0)|+\sum_{n=1}^\infty |a_n|r^n + \sum_{n=1}^\infty |b_n|r^n \le 1~\mbox{ for all $\ds r \le \frac{2}{4k+6}$}
$$
which for $k=0$ gives the classical Bohr's inequality and for $k=1$, this inequality contains the Bohr inequality for
sense-preserving harmonic mapping  $f(z)= h(z) + \overline{g(z)}$ of the disk $\ID$ with the Bohr radius $1/5$
(see \cite{bib:15}).
\er

\bthm\label{AKP2-th5}
Assume the hypotheses of Theorem \ref{AKP2-th4}. Then we have
$$
E_f(z):=
|h(z^m)|+|z^m|\,|h'(z^m)|+\sum_{n=2}^\infty |a_n|r^n + \sum_{n=1}^\infty |b_n|r^n \le 1 ~\mbox{ for all $r \le R^k_{m,2}$},
$$
where $R^k_{m,2}$ is the maximal positive root of the equation $\Lambda_m (r)=0$ with
$$
\Lambda_m (r)=
(1-r)(r^{2m}+2r^m-1)+2r(r+k)(1+r^m)^2
$$
and the constant $R^k_{m,2}$ cannot be improved.
\ethm
\begin{table}[!htbp] \begin{center}
\begin{tabular}{|l|l|}
\hline
 {\bf $m$}   & $R^1_{m,2}$ \\
  \hline
1  &  0.1671  \\
  \hline
2  &  0.240751  \\
  \hline
3  & 0.267472 \\
  \hline
4  & 0.276691\\
  \hline
5  & 0.279585  \\
  \hline
\end{tabular}
\end{center}
\caption{$R^1_{m,2}$ is the maximal positive root of the equation
$(1-r)(r^{2m}+2r^m-1)+2r(r+1)(1+r^m)^2=0$ \label{AKP2-th5-T5}}
\end{table}
\bpf
As in the proofs of Theorem \ref{AKP2-th4} and earlier theorems, we easily have
\begin{align*}
E_f(z) &\le \frac{r^m+a}{1+ar^m} +
\frac{r^m}{1-r^{2m}}\left[1-\left(\frac{r^m+a}{1+ar^m}\right)^2\right]+
(1-a^2)\frac{r^2}{1-r} + k(1-a^2)\frac{r}{1-r}\\
& =  1-\frac{(1-a)(1-r^m)}{1+ar^m} +  \frac{(1-a^2)r^m}{(1+ar^m)^2} +\frac{(1-a^2)(r+k)r}{1-r}\\
&= 1+\frac{(1-a)\Lambda_m(a,r)}{(1+ar^m)^2(1-r)},
\end{align*}
where
\begin{eqnarray*}
\Lambda_m(a,r) &=&  - (1-r^m)(1-r)(1+ar^m) +  r^{m}(1+a)(1-r) + (1+a)r(r+k)(1+ar^m)^2 \\
&=& (1-r)(ar^{2m}+2r^m-1) + r(1+a)(r+k)(1+ar^m)^2\le \Lambda_m(1,r) 
=\Lambda_m(r).
\end{eqnarray*}
The first inequality above is justified with the same reasoning as in the proofs of earlier theorems.

Now, we see that $E_f(z)\leq 1$ whenever $\Lambda_m(r) \le0$, which holds for $r\le R^k_{m,2}$,
where $R^k_{m,2}$ is the maximal positive root of the equation $\Lambda_m(r)=0$.

To show the sharpness of the radius $R^k_{m,2}$,  consider the function $f$ defined by \eqref{AKP-eq18}
with $\lambda \in (0,1]$. For this function, the corresponding expression for $E_f(z)$ with $z=r$ turned out to be
\begin{equation}\label{AKP-eq13}
E_f(z)= \frac{a+2r^m+ar^{2m}}{(1+ar^m)^2}+\frac{ar(1-a^2)(r+\lambda)}{1-ar}.
\end{equation}
The last expression is larger than $1$ if and only if $P^k_m(a,r)>0$, where
\begin{equation}\label{AKP-eq14}
P^k_m(a,r)=(1-ar)(ar^{2m}+2r^m-1) + ar(1+a)(r+\lambda)(1+ar^m)^2.
\end{equation}
By a simple calculation, we find that $P^k_m(a,r)$ is an increasing function of $a\in [0,1)$, and for each $r\in [0,1)$, so that
$$P^k_m(a,r)\le P^k_m(1,r) = (1-r)(r^{2m}+2r^m-1) + 2r(r+\lambda)(1+r^m)^2.
$$
Therefore, the expression (\ref{AKP-eq13}) is smaller than or equal to $1$ for all $a\in [0,1)$, only in the case
when $r\le R^{k}_{m,2}$ ($\lambda =k$). Finally, it also suggests that $a\to 1$ in (\ref{AKP-eq14}) shows that  the expression
(\ref{AKP-eq13}) is larger than $1$ if $r > R^k_{m,2}$. This completes the proof.
\epf

\br In Table \ref{AKP2-th5-T5}, we listed the values of $R^k_{m,2}$ for $k=1$ and for certain values of $m$.
If we allow $m\to\infty$ in Theorem \ref{AKP2-th5}, we obtain that
$$R^k_{m,2}\to
R^k_{2}:=\frac14\left(\sqrt{(2k+1)^2+8}-(2k+1)\right),
$$
where $R^k_{2}$ is the positive root of the equation $2x(x+k)+x-1=0$
and the conclusion of Theorem~\ref{AKP2-th5} takes the following form:
$$ |h(0)|+\sum_{n=2}^\infty |a_n|r^n + \sum_{n=1}^\infty |b_n|r^n \le 1 ~\text{ for all }~
r \le \frac14\left(\sqrt{(2k+1)^2+8}-(2k+1)\right).
$$
\er

\subsection*{Acknowledgements}
The work of S. Alkhaleefah and I. Kayumov is supported by the Russian Science Foundation under grant 18-11-00115.
The  work of the third author is supported by Mathematical Research Impact Centric Support of Department of 
Science \& Technology, India  (MTR/2017/000367).


\end{document}